OpenStudy (anonymous):

Write the equation -2x + 6y = 7 in polar form. How do I do this? Please help!

OpenStudy (anonymous):

oh no that is wrong sorry that would be a line through the origin, and this one is not

OpenStudy (anonymous):

let me see if i can do it with pencil and paper, might take a few

OpenStudy (anonymous):

does \(P=r\cos(\theta -\phi)\) look familiar?

OpenStudy (anonymous):

Yes, polar form.

OpenStudy (anonymous):

so what i think you need is \(\phi\) and \(P\)

OpenStudy (anonymous):

\[-2x + 6y = 7 \\ -2x+6y-7=0\]

OpenStudy (anonymous):

correct me if i am wrong but i think you need \(\sqrt{2^2+6^2}=\sqrt{40}=2\sqrt{10}\) and then divide all by that

OpenStudy (anonymous):

that gives \[-\frac{x}{\sqrt{10}}+\frac{3}{\sqrt{10}}-\frac{7}{\sqrt{10}}=0\]

OpenStudy (anonymous):

i might be off by a minus sign there, not sure

OpenStudy (anonymous):

no, i think that might be right

OpenStudy (anonymous):

That's what I did, but the answer I got wasn't one of my options. My options are: \[7\sqrt10=r \cos(\theta-108°)\] \[\sqrt10/20=r \cos(\theta-108°)\] \[7\sqrt10/20=r \sin(\theta-108°)\] \[7\sqrt10/20=r \cos(\theta-108°)\]

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

evidently the angle is \(108\) because that is all your choices here we have \(\tan(\phi)=3\) so \(\phi =\tan^{-1}(3)\) lets check that

OpenStudy (anonymous):

ah damn in the wrong quadrant \(180-71.57=108\) i guess if you round to the nearest degree

OpenStudy (anonymous):

you sure it is not \(\frac{7}{\sqrt{10}}\) out front?

OpenStudy (anonymous):

Yes, I got that too, but it's the value of P that I don't understand.

OpenStudy (anonymous):

Yeah, I'm sure

OpenStudy (anonymous):

oooooh i made a mistake i see it now

OpenStudy (anonymous):

you divide by \(\sqrt{40}=2\sqrt{10}\) i cancelled the 2 for the x and y term, but also cancelled one for \(-7\) which was a mistake it should be \(p=\frac{7}{2\sqrt{10}}\)

OpenStudy (anonymous):

That's still not one of my options though :(

OpenStudy (anonymous):

rationalize the denominator and get \[\frac{7}{2\sqrt{10}}=\frac{7\sqrt{10}}{20}\]

OpenStudy (anonymous):

now it is one of your options right?

OpenStudy (anonymous):

Ohh I was plugging it into wolfram alpha wrong ahaha yes thank you so much!!

OpenStudy (anonymous):

last option i believe is the right one